A dog walking to the right at 1.5 m/s speeds up to chase a cat with a constant acceleration of 12 m/s2.
What is the velocity of the dog after speeding up for 3 m?
1
Expert's answer
2019-11-19T07:46:26-0500

We have:


"V^2-v^2=2as"

"V^2-1.5^2=2(12)(3)"

"V=8.6\frac{m}{s}"